OpenStudy (anonymous):

The solution to the the antiderivative problem Sf(x) dx turned out to be sinx 􀀀- xcos x + C. What is f(x)?

OpenStudy (anonymous):

used the derivative with respect to x of that function to come up with -xsinx

OpenStudy (anonymous):

oops, looks like it's supposed to be xsinx...is that right?

OpenStudy (welshfella):

you need to do the reverse - that is find the derivative of sin x - x cosx

OpenStudy (welshfella):

you can find the derivative of x cos x by using the Product Rule

jimthompson5910 (jim_thompson5910):

cleetus779, you are correct if y = sin(x) - x*cos(x), then dy/dx = x*sin(x) so that's why \[\Large \int x\sin(x)dx = \sin(x) - x\cos(x)\]

jimthompson5910 (jim_thompson5910):

oops plus C of course

OpenStudy (welshfella):

d(x cosx) dx = x * -sinx + cos x

OpenStudy (anonymous):

thanks for checking guys

OpenStudy (welshfella):

yw